A contractor is laying out the sides of a rectangular porch. She knows that the porch will form a rectangle if it can be broken up into two right triangles, as shown below.

The lengths of the two longer sides of the porch are 24 feet and the lengths of the two shorter sides are 17 feet. Enter the length of the diagonal of the porch in feet, rounded to the nearest tenth.

Answer by Boreal(15235)

24^2+17^2=diagnoal^2
576+289=865=diagonal ^2
diagonal=sqrt(865)=29.4 feet